OpenStudy (anonymous):

Giving medel!!! Find f(2) if f(x) = (x + 1)^2.

OpenStudy (anonymous):

just substitute x by 2 and perfomr the operations

OpenStudy (nurali):

f(x) = (x + 1)^2 f(2)=(2+1)^2 f(2)=3^2 f(2)=9